PLEASE HELP! Dealing with molality and pressure

Calculate the morality of a solution containing 0.
Answer #1

16)
molality = number of moles of solute / mass of solvenet in Kg
                  = 0.755/1.75
                  = 0.=431 m

17)
molality = number of moles of solute / mass of solvenet in Kg
0.66 = 0.15 / mass of solvenet in Kg
mass of solvenet in Kg = 0.227 Kg
mass of solvent= 227 g
mass of solute = number of moles * molar mass
                               = 0.15*80.04
                               = 12
mass of solution = 227 g + 12 g = 239 g
Answer: b

I am allowed to answer only 1 question at a time